. Which of the following statements are true (T) and which are false (F)?
 (b) 400 is the predecessor of 399.
(c) Zero is the smallest whole number
(d) 600 is the successor of 599.
 

Grade:12th pass

3 Answers

Harshit Singh
2 years ago
Dear Student



(b) 400 is the predecessor of 399.
Solution: False
The predecessor of 399 is 398 Since, (399 – 1 = 398)

(c) Zero is the smallest whole number.
Solution: True
Zero is the smallest whole number

(d) 600 is the successor of 599.
Solution: True
Since (599 + 1 = 600)



Thanks
Vijay
2 years ago
Dear friends -
 
Statement 1 – 400 is predecessor of 399.
answer – False 
                Explanation – Predecessor means comes before the given number . 
                                         Here predecessor of 399 is 398.
 
Statement 2 – Zero is the smallest whole number.
Answer – True 
                 Explanation – Whole number are a set of number whuch includes all positive integers including zero.
 
Statement 3 – 600 is successor of 599
Answer – True 
                Explanation – successor means simply the next one . so 600 is the successor of 599 ( 599 +1 ).
 
Thanks
Debarati
2 years ago
b) False
Predecessor of 399= 399 – 1= 388
c) True
Whole no. starts from 0 and continues till infinity
d) True
Successor of 599= 599 + 1= 600